OpenStudy (anonymous):

Solve the following system of equations for Y: 5x-2y =10 25x^2 +4y^2 =100

OpenStudy (misty1212):

if \[5x-2y=10\] then \[5x=10+2y\]

OpenStudy (misty1212):

since \(25x^2=(5x)^2\) you can solve \[(10+2y)^2+4y^2=100\]

OpenStudy (nincompoop):

http://prntscr.com/79gdki